OpenStudy (anonymous):

How many square metres of canvas is required for a conical tent whose height is 3.5m and radius of whose base is 12m? (Use pie= 22/7)

OpenStudy (anonymous):

Assuming you don't want to cover the base in canvas, you'd subtract the area of the circlular base from the total surface area of the cone

OpenStudy (anonymous):

Surface area of a cone is given by pi(r^2) + pi (r)(l) 22/7 * 12^2 + 22/7 (12)(3.5) surface area is 207 squaremeters. if u just want the top part then its just this pi (r)(l)

OpenStudy (amistre64):



OpenStudy (amistre64):



OpenStudy (amistre64):

C = 2pir; = 24pi rn (new radius) = sqrt(3.5^2+12^2) = sqrt(...hmmm); sqrt(156.25) 35 35 925 150 150 ---- 12.25 144 ------ 156.25 ? C = t rn area = t rn^2/ 2 area = C rn^2/2 rn = C rn/2 maybe :)

OpenStudy (amistre64):

let me dbl chk that :) part of the circumference of flattened cone with a slit is the circumference of the cone base: 24 pi radius of flattened cone is slant height of cone; sqrt((3.5)^2 + 12^2) = 12.5 area of a sector: angle*r^2 / 2 arclength (C) of a sector is: angle * r angle = C/r C r^2 ----- = Cr / 2 2 r 12.5 * 24pi / 2 = 12.5 * 12 * 22/7 = abt 471 ; so i might be off somewhere :)

OpenStudy (amistre64):

yeah, so that should be good :)

OpenStudy (anonymous):

it is 471..